A and B have $ 18,000 between them. If A spends 1/3 of his share, and B spends 1/4 of his share, they will have $ 13,000 left. How much did each have at the beginning? show the solution.
Answer by richwmiller(17219)

a+b=18000
a-a/3+b-b/4=13000
12/12*a-12/12*a/3+12/12*b-12/12*b/4=13000
12/12*a-4/12*a+12/12*b-3/12*b=13000
8/12*a+9/12*b=13000
2/3*a+3/4*b=13000
a+b=18000
a = 6000, b = 12000